# Critical behavior of a programmable time-crystal lattice gas

### v1

@@ I Introduction
In quantum systems, continuous time crystals are forbidden in equilibrium short-range systems due to various no-go theorems Bruno (2013); Nozières (2013); Watanabe and Oshikawa (2015); Kozin and Kyriienko (2019). These limitations are overcome in non-equilibrium dissipative settings, which allow for the formation of continuous time crystals Iemini et al. (2018); Buča et al. (2019); Keßler et al. (2019); Carollo et al. (2020); Carollo and Lesanovsky (2022). Such a continuous time crystal was recently observed experimentally in an atom-cavity system Kongkhambut et al. (2022). Discrete time crystals, on the other hand, can appear as a subharmonic response to periodic (Floquet) driving. These have been theoretically predicted Else et al. (2016); Moessner and Sondhi (2017); Yao et al. (2017); Gong et al. (2018); Gambetta et al. (2019a); Khemani et al. (2019); Lazarides et al. (2020); Else et al. (2020) and experimentally observed in isolated Zhang et al. (2017); Choi et al. (2017); Rovny et al. (2018); Smits et al. (2018); Autti et al. (2018); O’Sullivan et al. (2020); Kyprianidis et al. (2021); Randall et al. (2021); Xiao et al. (2022) and dissipative Keßler et al. (2021); Kongkhambut et al. (2021) systems.
Classical systems can also exhibit time-crystalline order. For instance, discrete time crystal phases have been predicted in a periodically driven two-dimensional ($2d$) Ising model Gambetta et al. (2019b) and in a one-dimensional ($1d$) system of coupled nonlinear pendula at finite temperature Yao et al. (2020a), with experimental demonstrations in a classical network of dissipative parametric resonators Heugel et al. (2019). A continuous time crystal has been observed in a classical $2d$ array of plasmonic metamolecules, displaying superradiant-like transmissivity oscillations Liu et al. (2023). Despite these advances, a comprehensive approach to creating programmable time-crystal phases remains elusive.
@@
[…]
Mathematically, the action of the packing field can be seen as a controlled excitation of the first Fourier mode of the density field around the instantaneous center-of-mass position Hurtado-Gutiérrez et al. (2020). A natural question then concerns the system response to the excitation of higher, $m$th-order modes (with $m>1$). We have recently demonstrated using hydrodynamics and a local stability analysis Hurtado-Gutiérrez et al. (2025) that this excitation mechanism opens the door to fully programmable continuous time-crystal phases in driven diffusive fluids, characterized by an arbitrary number $m$ of rotating condensates, see Fig. 1, which can be further enhanced with higher-order modes. In particular, hydrodynamics allows the derivation of general properties of the condensates density profiles and velocities, as well as a scaling property of higher-order traveling condensates in terms of first-order ones. These findings were illustrated in Hurtado-Gutiérrez et al. (2025) for several paradigmatic driven diffusive systems Spohn (2012); Kipnis et al. (1982); Spitzer (1970); Derrida (1998); Katz et al. (1984); Hager et al. (2001); Krapivsky (2013), where programmable time-crystal phases were demonstrated and characterized in all cases, finding along the way a novel explosive time-crystal phase transition for certain nonlinear transport coefficients. Programmable in this context means that we can control on demand the number, shape, and velocity of the emerging condensates. In this way, these results demonstrate the utility and rich possibilities of this promising route to time crystals, which may serve as a robust platform to engineer these complex nonequilibrium phases of matter and study how to exploit them at the technological level.
@@ II The packing-field route to time crystals
We hence consider a stochastic lattice gas with $N$ particles moving on a one-dimensional ($1d$) periodic ring with $L\geq N$ lattice sites, so that the global density is $\rho_{0}\equiv N/L\leq 1$. Each lattice site $k\in[1,L]$ might be either empty (occupation number $n_{k}=0$) or occupied by one particle at most ($n_{k}=1$), defining a system configuration $\mathbf{n}\equiv\{n_{k},k\in[1,L]\}$. A particle at site $k$ can jump randomly to a empty neighboring site in the clockwise (--) or anti-clockwise ($+$) direction with rates $p^{\pm}_{k}(\mathbf{n})=\frac{1}{2}\,\mathrm{e}^{\pm E_{k}(\mathbf{n})/L}$, with $E_{k}(\mathbf{n})$ an external field that may depend on the particular site $k$ and the corresponding system configuration $\mathbf{n}$. Following the ideas introduced in Hurtado-Gutiérrez et al. (2020, 2025) and discussed above, the TCLG is characterized by an external field with two components, $E_{k}(\mathbf{n})=\epsilon+\lambda\mathcal{E}^{(m)}_{k}(\mathbf{n})$. Here $\epsilon$ is a constant driving field pushing the particles homogeneously in a given direction, while $\lambda\geq 0$ is the coupling to an $m$th-order packing field $\mathcal{E}^{(m)}_{k}(\mathbf{n})$ defined as
@@
[…]
The packing field mechanism works by restraining the motion of particles ahead of the closest localization center, i.e. for nearby lattice sites $k$ where $\mathcal{E}^{(m)}_{k}(\mathbf{n})<0$, see Eq. (1), while pushing particles lagging behind this point (where $\mathcal{E}^{(m)}_{k}(\mathbf{n})>0$), see Fig. 1. The strength of the packing mechanism is proportional to the coupling constant $\lambda$ and the magnitude $\lvert z_{m}\rvert$ of the packing order parameter, which measures the concentration of particles around the emergent localization centers. For large enough values of $\lambda$, the packing field leads to a nonlinear feedback mechanism which amplifies the density fluctuations naturally present in the system, resulting eventually in a phase transition to a time-crystal phase for large enough values of $\lambda$, exhibiting the fingerprints of spontaneous time-translation symmetry breaking Hurtado-Gutiérrez et al. (2020, 2025). Interestingly, the packing field can also be expressed as a long-range pairwise interaction between all particles. Indeed, by expanding the value of $z_{m}$, and using some basic trigonometric identities, it is easy to show that Hurtado-Gutiérrez et al. (2020, 2025)
@@
The evolution of the TCLG model here introduced is described at the microscopic level by a Markovian master equation for the probability of a configuration $\mathbf{n}$ at time $t$ Hurtado-Gutiérrez et al. (2020, 2025). In the macroscopic limit $L\to\infty$, using a local equilibrium ansatz and performing a diffusive scaling of space ($x=k/L\in[0,1]$) and time ($t=\tau/L^{2}$, with $\tau$ being the microscopic time), one can show that the model is characterized by a density field $\rho(x,t)$, with $0\leq\rho\leq 1$, which evolves according to the following hydrodynamic equation Spohn (2012); Derrida ; Bertini et al. (2015); Lasanta et al. (2015); Gutiérrez-Ariza and Hurtado (2019)
@@
with periodic boundary conditions, a diffusion coefficient $D(\rho)=1/2$, and a mobility $\sigma(\rho)=\rho(1-\rho)$. The external field takes the expected form $E_{x}[\rho]=\epsilon+\lambda\mathcal{E}_{x}^{(m)}[\rho]$, with a constant driving field $\epsilon$ and a macroscopic packing field now given by $\mathcal{E}_{x}^{(m)}[\rho]=\lvert z_{m}\rvert\sin(\varphi_{m}-2\pi xm)$ and written in terms of the macroscopic $m$th-order packing order parameter
[…]
@@
$\lambda_{c}^{(m)}=4\pi m\frac{D(\rho_{0})\rho_{0}}{\sigma(\rho_{0})}\,.$ (6)
Replacing the values of $\sigma(\rho_{0})$ and $D(\rho_{0})$ in the TCLG model we hence obtain
$\lambda_{c}^{(m)}=\frac{2\pi m}{1-\rho_{0}}\,.$ (7)
@@
[…]
@@ III Monte Carlo simulations of the time crystal lattice gas
We proceed now to verify the forecasted instability and the different hydrodynamic predictions in Monte Carlo simulations of the microscopic TCLG model introduced in Section §II, comparing with the theoretical results obtained from the numerical integration of the hydrodynamic evolution equations whenever possible. In particular, we carried out extensive discrete-time Monte Carlo simulations Bortz et al. (1975); Gillespie (1977); Newman and Barkema (1999); Binder and Heermann (2010) at a constant global density $\rho_{0}=1/3$ and varying values of $L\in[480,2880]$, so as to perform a detailed finite-size scaling analysis of the phase transition. Unless otherwise specified, we will work with a fixed driving field $\epsilon_{m}=2.5m$ for each packing order $m$. This scaling of the driving field with $m$ will allow us to compare condensate profiles for different values of $m$, as predicted by the hydrodynamic theory. We will however permit variations in $\epsilon_{m}$ later on in this section, in order to explore the effect of the driving field on the resulting particle condensates.
@@
Figure 3: Characterizing the phase transition. Monte Carlo results for the order parameter $\langle|z_{m}|\rangle$ [panels (a)-(c)], its scaled fluctuations $\langle\Delta|z_{m}|^{2}\rangle=L(\langle|z_{m}|^{2}\rangle-\langle|z_{m}|\rangle^{2})$ [panels (d)-(f)], and the Binder cumulant $U_{4}$ [panels (g)-(i)], for the TCLG model as a function of the packing field coupling $\lambda$ and measured for packing orders $m=1$ (left column), $m=2$ (central column), and $m=3$ (right column), global density $\rho_{0}=1/3$, driving field $\epsilon_{m}=2.5m$, and different values of $L$ (see legend). The red dashed lines in panels (a)-(c) correspond to the hydrodynamic theory prediction for $\langle|z_{m}|\rangle(\lambda)$, while the vertical dashed lines in all panels signal the predicted critical point $\lambda_{c}^{(m)}$, see Eq. (7).
@@
The phase transition is most evident at the configurational level, so we measured the average particle density profiles along the $1d$ ring for different values of the packing field coupling $\lambda$ across the predicted critical point $\lambda_{c}^{(m)}$, see Eq. (7), different packing orders $m=1,~{}2,~{}3$, and varying $L$, see Fig. 2. Due to the system periodicity, and in order not to blur away the possible structure of the density field, we performed profile averages only after shifting the argument of the instantaneous packing order parameter to $x=1/2$. Note that this procedure leads to a spurious weak structure even in the homogeneous density phase ($\lambda<\lambda_{c}^{(m)}$) due to the fluctuations of the particles’ packing Hurtado and Garrido (2011), see Figs. 2.(a)-(c), equivalent to averaging noisy homogeneous density profiles around their (random) center of mass. On the other hand, supercritical ($\lambda>\lambda_{c}^{(m)}$) density fields exhibit a much more pronounced structure resulting from the appearance of traveling particle condensates. In this way, Figs. 2.(a)-(c) thus confirm that before the critical point ($\lambda<\lambda_{c}^{(m)}$), density profiles remain homogeneous despite the system sustains a net particle current in the direction of the driving field. However, as the packing field coupling $\lambda$ is increased beyond the predicted critical value $\lambda_{c}^{(m)}$, see Eq. (7), a collection of $m$ equivalent particle condensates emerges. These condensates travel along the $1d$ lattice ring with an approximately constant speed, with weak velocity fluctuations vanishing in the $L\to\infty$ limit. The emergent matter waves become more prominent and compact as the coupling $\lambda$ increases, see Figs. 2.(g)-(i). Interestingly, the density profiles of the condensates are slightly asymmetric in shape, and this asymmetry increases and becomes more noticeable as $\lambda$ grows. We will further discuss this asymmetry below, and the role of the constant driving field in its control. Notice also that, in all cases studied, there is a remarkable agreement between the predictions of the hydrodynamic theory for the density profile and the simulation results of the TCLG, see Fig. 2, even for relatively small lattice sizes.
@@
Figure 4: Rigidity of the time-crystal phase. Panels (a)-(c) show the oscillations of the average density at a given lattice site, $\langle\rho(x_{0},t)\rangle$, for packing orders $m=1,~{}2,~{}3$, respectively, parameters $\epsilon_{m}=2.5m$ and $\lambda=2\lambda_{c}^{(m)}$, and for different lattice sizes $L=240,~{}480,~{}960,~{}1920,~{}2880$ (plotted in increasing color intensity). Note that the amplitude of oscillations is damped in time for finite lattice sizes. The black dashed line shows the hydrodynamic solution in each case, while the gray dash-dotted line displays the exponential envelope of the $L=240$ curves. Panel (d) displays the measured damping rate $\gamma_{L}$ of the oscillations as a function of $1/L$, showing how it vanishes in the macroscopic limit, a signature of the rigidity of the resulting time-crystal phases in this limit. Panel (e) illustrates the convergence of the measured (finite-size) angular speed $\omega_{m}$ of the traveling wave to the hydrodynamic prediction, shown as a black horizontal dashed line. In all cases, error bars are not shown because they are smaller than the markers. For all choices of the parameters, results were obtained from averaging at least 1000 runs of the simulations.
@@
Figure 5: Data collapses around the critical region. Scaling plots for the order parameter $\langle|z_{m}|\rangle$ (top row), the susceptibility $\langle\Delta|z_{m}|^{2}\rangle$ (middle row), and the Binder cumulant $U_{4}$ (bottom row), using all data shown in Fig. 3 (corresponding to $\rho_{0}=1/3$, $L\in[480,2880]$, $m=1,~{}2,~{}3$ and $\epsilon(m)=2.5m$) and the critical exponents $\nu=2$, $\gamma=1$, $\beta=1/2$ corresponding to the (mean field) Kuramoto universality class. Vertical dashed lines signal the critical point.

### v2

@@ I Introduction
In quantum systems, continuous time crystals are forbidden in equilibrium short-range systems due to various no-go theorems Bruno (2013); Nozières (2013); Watanabe and Oshikawa (2015); Kozin and Kyriienko (2019). These limitations are overcome in nonequilibrium dissipative settings, which allow for the formation of continuous time crystals Iemini et al. (2018); Buča et al. (2019); Keßler et al. (2019); Carollo et al. (2020); Carollo and Lesanovsky (2022). Such a continuous time crystal was recently observed experimentally in an atom-cavity system Kongkhambut et al. (2022). Discrete time crystals, on the other hand, can appear as a subharmonic response to periodic (Floquet) driving. These have been theoretically predicted Else et al. (2016); Moessner and Sondhi (2017); Yao et al. (2017); Gong et al. (2018); Gambetta et al. (2019a); Khemani et al. (2019); Pizzi et al. (2019); Lazarides et al. (2020); Else et al. (2020); Pizzi et al. (2021a); Yousefjani et al. (2025) and experimentally observed in isolated Zhang et al. (2017); Choi et al. (2017); Rovny et al. (2018); Smits et al. (2018); Autti et al. (2018); O’Sullivan et al. (2020); Kyprianidis et al. (2021); Randall et al. (2021); Xiao et al. (2022) and dissipative Keßler et al. (2021); Kongkhambut et al. (2021) systems.
Classical systems can also exhibit time-crystalline order Pizzi et al. (2021b, c). For instance, discrete time crystal phases have been predicted in a periodically driven two-dimensional ($2d$) Ising model Gambetta et al. (2019b) and in a one-dimensional ($1d$) system of coupled nonlinear pendula at finite temperature Yao et al. (2020a), with experimental demonstrations in a classical network of dissipative parametric resonators Heugel et al. (2019). A continuous time crystal has been observed in a classical $2d$ array of plasmonic metamolecules, displaying superradiant-like transmissivity oscillations Liu et al. (2023). Despite these advances, a comprehensive approach to creating programmable time-crystal phases remains elusive.
@@
Figure 1: The packing-field mechanism. (a) A stochastic particle fluid in the presence of a constant driving field $\epsilon$ sustains a net current of particles with a homogeneous density structure on average subject to small density fluctuations. (b) By switching on a $m$-th-order packing field $\mathcal{E}^{(m)}$ with a coupling $\lambda$ beyond a critical value $\lambda_{c}$, an instability is triggered to a complex time-crystal phase characterized by the emergence of $m$ rotating particle consensates with a velocity controlled by the driving field. (c) Sketch showing the working of the packing field $\mathcal{E}^{(m)}$ for different packing orders $m=1,2,3$. The top row shows possible condensate configurations for varying $m$, together with the corresponding packing order parameter $z_{m}=\lvert z_{m}\rvert e^{i\varphi_{m}}$ (orange vector). Its magnitude $\lvert z_{m}\rvert$ measures the amount of particle packing around the $m$ emergent localization centers located at angular positions $\phi_{m}^{(j)}=(\varphi_{m}+2\pi j)/m$ (see green arrows), with $j\in[0,m-1]$, i.e. the arguments of $(\sqrt[m]{z_{m}})_{j}=\sqrt[m]{z_{m}}\textrm{e}^{\textrm{i}2\pi j/m}$. The red, external arrows in top panels signal the local direction of the packing field, which hinders particle motion ahead of each localization point and pushes particles lagging behind, promoting in this way particle packing. The bottom panels display the value of the packing field as a function of the angular position along the ring, while the green dashed lines mark the emergent localization centers at $\phi_{m}^{(j)}$.
Mathematically, the action of the packing field can be seen as a controlled excitation of the first Fourier mode of the density field around the instantaneous center-of-mass position Hurtado-Gutiérrez et al. (2020). A natural question then concerns the system response to the excitation of higher, $m$th-order modes (with $m>1$). We have recently demonstrated using hydrodynamics and a local stability analysis Hurtado-Gutiérrez et al. (2025) that this excitation mechanism opens the door to fully programmable continuous time-crystal phases in driven diffusive fluids, characterized by an arbitrary number $m$ of rotating condensates, see Fig. 1, which can be further enhanced with higher-order modes. In particular, hydrodynamics allows the derivation of general properties of the condensates density profiles and velocities, as well as a scaling property of higher-order traveling condensates in terms of first-order ones. These findings were illustrated in Hurtado-Gutiérrez et al. (2025) for several paradigmatic driven diffusive systems Spohn (2012); Kipnis et al. (1982); Spitzer (1970); Derrida (1998); Katz et al. (1984); Hager et al. (2001); Krapivsky (2013), where programmable time-crystal phases were demonstrated and characterized in all cases, finding along the way a novel explosive time-crystal phase transition, i.e., a discontinuos transition to a time-crystal phase, for certain nonlinear transport coefficients. Programmable in this context means that we can control on demand the number, shape, and velocity of the emerging condensates. In this way, these results demonstrate the utility and rich possibilities of this promising route to time crystals, which may serve as a robust platform to engineer these complex nonequilibrium phases of matter and study how to exploit them at the technological level.
@@ II The packing-field route to time crystals
We hence consider a stochastic lattice gas with $N$ particles moving on a one-dimensional ($1d$) periodic ring with $L\geq N$ lattice sites, so that the global density is $\rho_{0}\equiv N/L\leq 1$. Each lattice site $k\in[1,L]$ might be either empty (occupation number $n_{k}=0$) or occupied by one particle at most ($n_{k}=1$), defining a system configuration $\mathbf{n}\equiv\{n_{k},k\in[1,L]\}$. A particle at site $k$ can jump randomly to a empty neighboring site in the clockwise (-) or anti-clockwise (++) direction with rates $p^{\pm}_{k}(\mathbf{n})=\frac{1}{2}\,\mathrm{e}^{\pm E_{k}(\mathbf{n})/L}$, with $E_{k}(\mathbf{n})$ an external field that may depend on the particular site $k$ and the corresponding system configuration $\mathbf{n}$. Following the ideas introduced in Hurtado-Gutiérrez et al. (2020, 2025) and discussed above, the TCLG is characterized by an external field with two components, $E_{k}(\mathbf{n})=\epsilon+\lambda\mathcal{E}^{(m)}_{k}(\mathbf{n})$. Here $\epsilon$ is a constant driving field pushing the particles homogeneously in a given direction, while $\lambda\geq 0$ is the coupling to an $m$th-order packing field $\mathcal{E}^{(m)}_{k}(\mathbf{n})$ defined as
@@
with magnitude $\lvert z_{m}\rvert$ and argument $\varphi_{m}$, which sets the angular position of the first condensate (given by $\varphi_{m}/m$), as we shall explain below. The constant driving $\epsilon$ gives rise to a net particle current in the desired direction, and as we will see below, controls the velocity of the resulting particle condensates and the asymmetry of the associated density waves. On the other hand, the packing field $\mathcal{E}^{(m)}_{k}(\mathbf{n})$ pushes particles locally towards $m$ emergent localization centers where particles are most clustered, with an amplitude proportional to the instantaneous magnitude of the packing order parameter $\lvert z_{m}\rvert$ and the coupling constant $\lambda$. The angular position along the $1d$ ring of the $m$ emergent localization centers is given by $\phi_{m}^{(j)}=(\varphi_{m}+2\pi j)/m$, with $j\in[0,m-1]$, i.e. the arguments of the $m$ roots of the complex $m$th-order packing order parameter, $(\sqrt[m]{z_{m}})_{j}=\sqrt[m]{z_{m}}\textrm{e}^{\textrm{i}2\pi j/m}$, as illustrated in Fig. 1.(c).
The packing field mechanism works by restraining the motion of particles ahead of the closest localization center, i.e. for nearby lattice sites $k$ where $\mathcal{E}^{(m)}_{k}(\mathbf{n})<0$, see Eq. (1), while pushing particles lagging behind this point (where $\mathcal{E}^{(m)}_{k}(\mathbf{n})>0$), see Fig. 1. The strength of the packing mechanism is proportional to the coupling constant $\lambda$ and the magnitude $\lvert z_{m}\rvert$ of the packing order parameter, which measures the concentration of particles around the emergent localization centers. For large enough values of $\lambda$, the packing field leads to a nonlinear feedback mechanism which amplifies the density fluctuations naturally present in the system, resulting eventually in a phase transition to a time-crystal phase, exhibiting the fingerprints of spontaneous time-translation symmetry breaking Hurtado-Gutiérrez et al. (2020, 2025). Interestingly, the packing field can also be expressed as a long-range pairwise interaction between all particles. Indeed, by expanding the value of $z_{m}$, and using some basic trigonometric identities, it is easy to show that Hurtado-Gutiérrez et al. (2020, 2025)
@@
The evolution of the TCLG model here introduced is described at the microscopic level by a Markovian master equation for the probability of a configuration $\mathbf{n}$ at time $t$ Hurtado-Gutiérrez et al. (2020, 2025). In the macroscopic limit $L\to\infty$, using a local equilibrium ansatz and performing a diffusive scaling of space ($x=k/L\in[0,1]$) and time ($t=\tau/L^{2}$, with $\tau$ being the microscopic time), one can show that the model is characterized by a density field $\rho(x,t)$, with $0\leq\rho\leq 1$, which evolves according to the usual hydrodynamic equation in one-dimensional driven-diffusive systems Spohn (2012); Derrida ; Bertini et al. (2015); Lasanta et al. (2015); Gutiérrez-Ariza and Hurtado (2019)
@@
with periodic boundary conditions. In the TCLG model, the diffusion coefficient is constant, $D(\rho)=1/2$, and the mobility is given by $\sigma(\rho)=\rho(1-\rho)$. As expected, the external field depends on the whole density profile and takes the form $E_{x}[\rho]=\epsilon+\lambda\mathcal{E}_{x}^{(m)}[\rho]$, with a constant driving field $\epsilon$ and a macroscopic packing field now given by $\mathcal{E}_{x}^{(m)}[\rho]=\lvert z_{m}\rvert\sin(\varphi_{m}-2\pi xm)$ and written in terms of the macroscopic $m$th-order packing order parameter
$z_{m}[\rho]=\frac{1}{\rho_{0}}\int_{0}^{1}dx\rho(x,t)\textrm{e}^{\textrm{i}2\pi mx}\equiv\lvert z_{m}\rvert\textrm{e}^{\textrm{i}\varphi_{m}}\,$, (5)
where $\rho_{0}=\int_{0}^{1}dx\,\rho(x,t)$ is the (constant) global density of the system.
@@
$\lambda_{c}^{(m)}=4\pi m\frac{D(\rho_{0})\rho_{0}}{\sigma(\rho_{0})}=\frac{2\pi m}{1-\rho_{0}}\,.$ (6)
@@
On the other hand, using the current field $j(x,t)\equiv-D(\rho)\partial_{x}\rho+\sigma(\rho)E_{x}[\rho]$, see Eq. (4), the excess of the average current $J=\tau^{-1}\int_{0}^{\tau}dt\int_{0}^{1}dx~j(x,t)$ with respect to the homogeneous-phase average current $J_{0}=\sigma(\rho_{0})\epsilon$ can be shown to be $J-J_{0}\propto\sigma^{\prime\prime}(\rho_{0})\epsilon$ right after the instability Hurtado-Gutiérrez et al. (2025). In this way the particle current will be larger or smaller than the homogeneous-phase flow depending on the mobility curvature for density $\rho_{0}$. This highlights the relevance of transport coefficients in the system’s response to the packing field, which enhances or lowers the current and the wave velocity depending on the mobility derivatives. For the particular case of the TCLG, where $\sigma^{\prime\prime}(\rho_{0})=-2<0$, we thus expect a negative excess current ($J<J_{0}$), compatible with the emergence of condensates that jam particle dynamics Pérez-Espigares et al. (2013); Hurtado-Gutiérrez et al. (2020, 2025).
@@ III Monte Carlo simulations of the time crystal lattice gas
We proceed now to verify the forecasted instability and the different hydrodynamic predictions in Monte Carlo simulations of the microscopic TCLG model introduced in Section §II, comparing with the theoretical results obtained from the numerical integration of the hydrodynamic evolution equation (4) whenever possible. In particular, we performed standard discrete-time random-sequential Monte Carlo simulations Bortz et al. (1975); Gillespie (1977); Newman and Barkema (1999); Binder and Heermann (2010) of the TCLG, where at each step a particle is chosen at random and attempts a biased hop to the right with probability $p_{k}^{+}(\mathbf{n})/(p_{k}^{+}(\mathbf{n})+p_{k}^{-}(\mathbf{n}))$ or to the left with probability $p_{k}^{-}(\mathbf{n})/(p_{k}^{+}(\mathbf{n})+p_{k}^{-}(\mathbf{n}))$ taking into account the exclusion rule. These simulations were carried out at a constant global density $\rho_{0}=1/3$ and varying values of $L\in[480,2880]$, so as to perform a detailed finite-size scaling analysis of the phase transition. Unless otherwise specified, we will work with a fixed driving field $\epsilon_{m}=2.5m$ for each packing order $m$. This scaling of the driving field with $m$ will allow us to compare condensate profiles for different values of $m$, as predicted by the hydrodynamic theory. We will however permit variations in $\epsilon_{m}$ later on in this section, in order to explore the effect of the driving field on the resulting particle condensates.
@@
Figure 3: Characterizing the phase transition. Monte Carlo results for the order parameter $\langle|z_{m}|\rangle$ [panels (a)-(c)], its scaled fluctuations $\langle\Delta|z_{m}|^{2}\rangle=L(\langle|z_{m}|^{2}\rangle-\langle|z_{m}|\rangle^{2})$ [panels (d)-(f)], and the Binder cumulant $U_{4}$ [panels (g)-(i)], for the TCLG model as a function of the packing field coupling $\lambda$ and measured for packing orders $m=1$ (left column), $m=2$ (central column), and $m=3$ (right column), global density $\rho_{0}=1/3$, driving field $\epsilon_{m}=2.5m$, and different values of $L$ (see legend). The red dashed lines in panels (a)-(c) correspond to the hydrodynamic theory prediction for $\langle|z_{m}|\rangle(\lambda)$, while the vertical dashed lines in all panels signal the predicted critical point $\lambda_{c}^{(m)}$, see Eq. (6). The solid curves correspond to steady-state time averages of the microscopic simulations.
@@
The phase transition is most evident at the configurational level, so we measured the average particle density profiles along the $1d$ ring for different values of the packing field coupling $\lambda$ across the predicted critical point $\lambda_{c}^{(m)}$, see Eq. (6), different packing orders m=1,2,3m=1,~2,~3, and varying $L$, see Fig. 2. Due to the system periodicity, and in order not to blur away the possible structure of the density field, we performed profile averages only after shifting the argument of the instantaneous packing order parameter to $x=1/2$. Note that this procedure leads to a spurious weak structure even in the homogeneous density phase ($\lambda<\lambda_{c}^{(m)}$) due to the fluctuations of the particles’ packing Hurtado and Garrido (2011), see Figs. 2.(a)-(c), equivalent to averaging noisy homogeneous density profiles around their (random) center of mass. On the other hand, supercritical ($\lambda>\lambda_{c}^{(m)}$) density fields exhibit a much more pronounced structure resulting from the appearance of traveling particle condensates. In this way, Figs. 2.(a)-(c) thus confirm that before the critical point ($\lambda<\lambda_{c}^{(m)}$), density profiles remain homogeneous despite the system sustains a net particle current in the direction of the driving field. However, as the packing field coupling $\lambda$ is increased beyond the predicted critical value $\lambda_{c}^{(m)}$, see Eq. (6), a collection of $m$ equivalent particle condensates emerges. These condensates travel along the $1d$ lattice ring with an approximately constant speed, with weak velocity fluctuations vanishing in the $L\to\infty$ limit. The emergent matter waves become more prominent and compact as the coupling $\lambda$ increases, see Figs. 2.(g)-(i). Interestingly, the density profiles of the condensates are slightly asymmetric in shape, and this asymmetry increases and becomes more noticeable as $\lambda$ grows. We will further discuss this asymmetry below, and the role of the constant driving field in its control. Notice also that, in all cases studied, there is a remarkable agreement between the predictions of the hydrodynamic theory for the density profile and the simulation results of the TCLG, see Fig. 2, even for relatively small lattice sizes.
@@
Figure 4: Rigidity of the time-crystal phase. Panels (a)-(c) show the oscillations of the average density at a given lattice site, $\langle\rho(x_{0},t)\rangle$, for packing orders m=1,2,3m=1,~2,~3, respectively, parameters $\epsilon_{m}=2.5m$ and $\lambda=2\lambda_{c}^{(m)}$, and for different lattice sizes L=240,480,960,1920,2880L=240,~480,~960,~1920,~2880 (plotted in increasing color intensity). Note that the amplitude of oscillations is damped in time for finite lattice sizes. The black dashed line shows the hydrodynamic solution in each case, while the gray dash-dotted line displays the exponential envelope of the $L=240$ curves. Panel (d) displays the measured damping rate $\gamma_{L}$ of the oscillations as a function of $1/L$, showing how it vanishes in the macroscopic limit, a signature of the rigidity of the resulting time-crystal phases in this limit. Panel (e) illustrates the convergence of the measured (finite-size) angular speed $\omega_{m}$ of the traveling wave to the hydrodynamic prediction, shown as a black horizontal dashed line. In all cases, error bars are not shown because they are smaller than the markers. For all choices of the parameters, results were obtained from averaging at least 1000 runs of the simulations.
@@
Figure 5: Data collapses around the critical region. Scaling plots for the order parameter $\langle|z_{m}|\rangle$ (top row), the susceptibility $\langle\Delta|z_{m}|^{2}\rangle$ (middle row), and the Binder cumulant $U_{4}$ (bottom row), using all data shown in Fig. 3 (corresponding to $\rho_{0}=1/3$, $L\in[480,2880]$, m=1,2,3m=1,~2,~3 and $\epsilon(m)=2.5m$) and the critical exponents $\nu=2$, $\gamma=1$, $\beta=1/2$ corresponding to the (mean field) Kuramoto universality class. Vertical dashed lines signal the critical point.